Station Messages
15 Aug 2020 03:01 UTC 2020228+03:01 UTC

Overview Datums Bench Marks Data Query Site Visit Daily QC Messages Attributes

132: Bahia Grande Water Quality Station #1

26° 1' 51" N   97° 17' 48" W
BGWQ1
University of Texas at Brownsville and Texas Southmost College University of Texas at Brownsville and Texas Southmost College

2010090+1254 (acknowledged, tussery)

No data since 2010084.

2010266+0955 (retired, ussery)
2010095+1049 (retired, ussery)

Data sporadic.

2010060+1357 (retired, ussery)

No data since 2010053.

2010049+1358 (retired, ussery)

No data since 2010047+0600.

2009312+1226 (retired, jrizzo)

No data since 2009261. Data resumed.

2009236+1257 (retired, ussery)

No data received since 2009225+1700.

2009125+1305 (retired, ussery)

No data since 2009115.

2009120+1052 (retired, jrizzo)

No data since 079

2009076+1124 (retired, jrizzo)

Little data since 065. Data resumed.

2008233+1641 (retired, ussery)

No data since 2008230. No response from tower.

2008210+1312 (retired, ussery)

No data since 2008206+1800. Power outage at Los Fresnos.

2008200+1307 (retired, ussery)

No data since 2008112+1830.

2008113+1325 (retired, ussery)

No data since 2008112+1830.

2008098+1225 (retired, jrizzo)

No data since 2008061. Data resumed.

2008061+0519 (retired, jadams)
all : 2/29/2008 (02/29/2008 00:00 - 03/01/2008 00:00) :

Hydrolab connector had water in it. purged the water with silicone grease.

2007354+1305 (retired, jrizzo)

No data since 07345. Data resumed.

2007345+1204 (retired, jrizzo)

No data from station since 2007336. Data resumed.

2007317+1400 (retired, ussery)
all :

No data since 2007279. Hydrolabs sent in for repair.

2007249+1218 (retired, jrizzo)

Specific conductance at zero. SPCON resumed.

2007227+1347 (retired, ussery)

Freewave connects, no response from hydrolab. Bad hydrolab cable.

2007168+1025 (retired, jrizzo)

Little data since 2007141

2007135+1038 (retired, jrizzo)

No data received since 2007122+1230. Battery voltage at 9.

2007121+1301 (retired, ussery)

Comm problems, data spotty. Data consistent.

2007101+1116 (retired, jrizzo)

No data since 097. Data resumed

2007093+1132 (retired, jrizzo)

data incomplete since 2007091. Data recovery resumed.

2007079+1939 (retired, jdavis)

Minimal Data since 2006364 -- station removed from poll until it can be repaired -- stationed repaired

2006348+1144 (retired, jrizzo)

Data resumed briefly 2006340, no data since 2006341. O&M Thursday. Data has resumed

2006341+1413 (retired, ussery)

No data since 2006334. Data now being received.

2006321+1341 (retired, ussery)

No data since 2006313. Data now coming in.

2006269+1244 (retired, jrizzo)

Little data since 2006267 for all BGWQ stations indicates polling problem. JDavis resolved polling issue

2006227+2216 (retired, jdavis)

Sporadic data since 2006226+1700.

2006221+1313 (retired, ussery)

Data received but not imported for all BGWQs since 2006219+1900. Data is imported.

2006094+1052 (retired, jrizzo)

DO sensor fouled

2006093+1655 (retired, jadams)
do sat : 2006092 (04/02/2006 00:00 - 04/03/2006 00:00) :

sensor replaced

2006058+1234 (retired, jrizzo)
bat : 2.17.2006 (02/17/2006 00:00 - 02/18/2006 00:00) :

UT-Brownsville partners will make repairs to the station 2/18, replacement parts have been shipped to UT Brownsville.

2006058+1234 (retired, jrizzo)

Battery @ 7V, no data. Battery up, data coming in.

2006034+1757 (retired, jdavis)

Los Fresnos tower down.
PC is back up and collecting data. Satellite modem is out. Data is being downloaded automatically over telephone modem every 8 hours until satellite service is restored.
----
Satellite modem is back up and data is being transferred.

2006030+1248 (retired, jrizzo)

No data from station 132 or 134 - trouble connecting to tower. Will visit 1/24. Tower system replaced

2006011+2345 (retired, kjenkin)

No data since 2006002. Hydrolab cable needs replacing

2006006+2350 (retired, kjenkin)

Connect fine to station - no data from hydrolab since 002. Hydrolab replaced

2005348+1847 (retired, kjenkin)

Water depth measurement increased 2.5M within .5 hours? Hydrolab setting corrected

2005218+1407 (retired, kjenkin)

No data since 2005215.